A stone of mass 1kg is thrown vertically downward from a rope at a velocity of 2m/s. At the moment it hits the soils, the stone has a kinetic energy of 200J. What is the height from which the stone is thrown?



1
Expert's answer
2021-07-27T05:11:01-0400

"E=\frac{mv^2}2+mgh,"

"h=\frac E{mg}-\frac{v^2}{2g}=19.8~m."